# MDPE: A Multimodal Deception Dataset with Personality and Emotional Characteristics

### v1

@@ 4.1 Data Preprocessing
For the visual modality, we first unify the raw video to 30 fps of the frame rate, and crop and align faces via DLib Toolkit [47]. Figure 2 shows some examples of real and deceptive faces. Then, we use visual encoders to extract frame-level , followed by average pooling to compress them to the video level. For the audio modality, we use FFmpeg to separate the audio from the video and unify the audio format to 16kHz and mono. For the textual modality, we first extract transcript using Paraformer [48], an open-source ASR toolkit. After data preprocessing, for each sample $x_{i}$, we extract acoustic features $f_{i}^{a}\in\mathbb{R}^{d_{a}}$, textual features $f_{i}^{l}\in\mathbb{R}^{d_{l}}$, and visual features $f_{i}^{v}\in\mathbb{R}^{d_{v}}$, where $\left\{d_{m}\right\}_{m\in\left\{a,l,v\right\}}$ is the feature dimension for each modality.
4.2 Feature extraction
Different features lead to distinct results. To guide feature selection, we evaluate the performance of different features under the same experimental setup.
Visual Modality: Compared with handcrafted features, deep features extracted from supervised models are useful for facial expression recognition [49]. CLIP [50] is a multimodal model based on contrastive learning, where training utilizes text and images to construct positive and negative sample pairs. Pre-training is conducted on a dataset comprising 400 million pairs, resulting in strong generalization capabilities. And Vision Transformer (ViT) [51] is a transformer encoder model , pre-trained in a supervised manner on a large dataset of images. Images are presented to the model in the form of sequences of fixed-size patches (resolution of 16x16) and undergo linear embedding.
Acoustic Modality: We extracted the handcrafted feature extended Geneva Minimalistic Acoustic Parameter Set (eGeMAPS) [52], which contains 88 acoustic parameters designed specially for speech emotional recognition tasks, covering spectral, cepstral, and prosodic features . And Wav2vec [53] demonstrates the power of learning robust representations solely from speech audio, fine-tuning on transcribed speech, surpassing the best semi-supervised methods. It masks speech inputs in the latent space and addresses a contrastive task defined on quantized latent representations. It has been widely applied to downstream speech tasks. HUBERT [54] utilizes offline clustering steps to provide aligned target labels for prediction losses. A key component of this approach is applying prediction losses only in masked regions, forcing the model to learn representations that combine acoustic and language models on contiguous inputs. To better distinguish between speakers, a sentence mixture training strategy WavLM [55] is proposed, allowing for the unsupervised creation and merging of additional overlapping sentences during the training process.
Text Modality: Among all language models, BERT [56] and its variants are widely utilized, and it uses the masked language model and next sentence prediction objectives to learn word embeddings. Here, we extract the sbert-chinese-general-v2 features, which is based on the bert-base-chinese version of the BERT model and is trained on the million-level semantic similarity data set SimCLUE. ChatGLM [57] is an open-source conversational language model supporting bilingual question answering in both Chinese and English, based on the General Language Model (GLM) architecture. It demonstrats exceptional contextual understanding and more efficient inference capabilities. Baichuan [58] is an open-source large-scale model with 13 billion parameters. It features a larger size, more extensive training data, and more efficient inference capabilities.
Figure 2: Some examples of truth and deception faces.
4.3 Model Structure
@@
$h_{i}^{m}=\text{ReLU}\left(f_{i}^{m}W_{m}^{h}+b_{m}^{h}\right),{m\in\left\{a,l,v\right\}}$ (1)
$\hat{y}_{i}=\text{softmax}\left(h_{i}^{m}W_{m}^{d}+b_{m}^{d}\right),m\in\left\{a,l,v\right\}$ (2)
@@
$h_{i}=\text{Concat}\left(h_{i}^{a},h_{i}^{l},h_{i}^{v}\right)$ (3)
$\alpha_{i}=\text{softmax}\left(h_{i}^{T}W_{\alpha}+b_{\alpha}\right)$ (4)
@@
Similarly, for personality and emotional expression features, we use concatenation for feature fusion. We directly use personality scale scores as personality features. For emotional features, we train an emotion recognition model first, input all emotional expression samples into the emotion recognition model, and take the last fully connected layer features for average pooling as the emotion expression feature.
@@
Table 2: Unimodal results on MDPE. "P" denotes the addition of personality features,"E" denotes the addition of emotional features
Feature Accuracy AUC with P with E with P and E Accuracy AUC Accuracy AUC Accuracy AUC VIT 60.30% 0.577 61.27% 0.582 61.43% 0.615 61.55% 0.620 CLIP-base 58.54% 0.602 59.17% 0.611 58.32% 0.606 59.11% 0.605 CLIP-large 57.30% 0.574 58.34% 0.582 56.97% 0.555 57.67% 0.579 eGeMAPS 55.86% 0.563 57.22% 0.588 56.22% 0.577 56.89% 0.585 HUBERT-base 58.13% 0.615 62.38% 0.651 59.35% 0.617 62.12% 0.641 HUBERT-large 60.80% 0.636 62.07% 0.646 60.34% 0.621 61.87% 0.644 Wav2vec2-base 58.75% 0.581 59.74% 0.603 59.99% 0.594 59.84% 0.599 Wav2vec2-large 60.10% 0.582 61.88% 0.617 59.32% 0.592 62.10% 0.634 WavLM-base 61.66% 0.609 60.82% 0.606 60.16% 0.595 60.92% 0.593 WavLM-large 57.82% 0.599 60.31% 0.583 58.02% 0.607 60.52% 0.611 Sentence-BERT 61.76% 0.639 62.34% 0.651 63.21% 0.641 63.34% 0.657 ChatGLM2-6B 60.73% 0.648 61.45% 0.659 61.45% 0.667 61.56% 0.676 Baichuan-13B 61.87% 0.649 62.90% 0.667 63.32% 0.675 63.74% 0.683
4.4 Implementation Details
We select the dimension of latent representations from $\left\{64,128,256\right\}$. During training, we use the Adam [59] optimizer and choose the learning rate from $\left\{10^{-3},10^{-4}\right\}$. We set the maximum number of epochs to 300 and the weight decay to $10^{-5}$. Dropout [60] is also employed, and we select the rate from $\left\{0.2,0.3,0.4,0.5\right\}$ to alleviate the over-fitting problem. We randomly select 5 answers (3 truths and 2 deceptions) from 24 answers in all samples as the validation set, and the remaining 19 answers as the training set. To mitigate randomness, we run each experiment five times and report the average result. And we choose the cross-entropy as loss function and the accuracy and AUC as the evaluation metric.
4.5 Experiment Results
Unimodal Results In this section, we establish the unimodal benchmark for MDPE and report results in Table 2. We hope this benchmark can provide guidance for feature selection and point the way to developing powerful feature extractors. For the visual modalities, VIT achieves better results than CLIP, possibly because VIT is trained on supervised datasets and can reveal more deceptive clues than CLIP features that use text as a supervisory signal. For the acoustic modality, the deep features outperform handcrafted features. This may be because egemas are features related to emotions, using a emotion-related handcrafted acoustic feature may limit performance. In contrast, deep features can capture more universal acoustic representations for deception detection. For the textual modality, we focus on textual encoders that support Chinese (large language models are generally trained on multilingual corpora containing Chinese). And the textual modality can achieve best performance than the visual and acoustic modalities, which indicates that our dataset textual reveals more deceptive clues than other modalities.
Of course, we have also incorporated personality and emotional features into deception detection. Models trained with personality features almost often exhibit superior performance. These findings demonstrate the importance of personality in deception detection tasks. And the addition of emotional features has also improved the performance of deception detection models, but the improvement is not comparable to personality features, even some results have decreased. This may be because personality traits are directly usable features, and emotional features are extracted by emotion recognition models. The quality of features is also influenced by the emotion recognition models. In the future, better methods for using emotional expression features can be explored. By incorporating both personality and emotional features, the deception detection model achieved the highest performance in unimodal results, These results show that personality and emotional expression characteristics are indeed important for deception detection tasks, and using them can truly achieve deception detection modeling based on individual differences. It is worth mentioning that among all the unimodal results, Baichuan always achieved the best results. It is the model with the largest number of parameters among the features we used, and it further demonstrates the potential of large language models in deception detection.
Multimodal Results In Table 3, we select several well-performing unimodal features and report their fusion results. Experimental results demonstrate that multimodal fusion consistently improves performance. The reason lies in the fact that deception cues can be conveyed through multiple modalities. The integration of multimodal information allows the model to better comprehend the video content and accurately detection deception. Firstly, almost all features have achieved performance improvements in multimodal feature fusion, but the fusion of visual and acoustic modalities has hardly improved or even declined. It indicates that the addition of text features make the performance of the model tend to stabilize. This is consistent with human judgment. When people judge whether others are lying, they tend to express the truth or falsehood of the content, because it difficult to judge from visual or acoustic features. In the future, further exploration and research should be conducted on deceptive clues in visual and acoustic features.
Similar to unimodal results, models trained with personality features typically exhibit excellent performance. The addition of emotional features makes the performance of the model unstable. Some features have been improved, while others have decreased. Finally, the best result always comes from the three modalities fusion, and the three modalities fusion with personality and emotional characteristics has achieved the best performance.
Table 3: Multimodal results on MDPE. We select several well-performing unimodal features and report their fusion results. Here, “V”, “A” and “T” represent the visual, acoustic and textual modalities, respectively.
V A T Accuracy AUC with P with E with P And E Accuracy AUC Accuracy AUC Accuracy AUC VIT HBB - 61.76% 0.628 61.53% 0.617 61.14% 0.616 62.68% 0.628 VIT WMB - 60.88% 0.611 60.22% 0.599 60.02% 0.599 60.72% 0.587 CLB HBB - 61.02% 0.618 60.14% 0.606 60.53% 0.614 61.45% 0.625 VIT - Bai 63.31% 0.664 63.47% 0.678 63.31% 0.676 63.48% 0.672 CLB - Bai 63.31% 0.666 63.52% 0.677 63.38% 0.667 63.10% 0.664 CLL - Bai 64.15% 0.665 63.94% 0.679 63.98% 0.672 64.04% 0.678 - HBL Bai 62.69% 0.658 63.42% 0.665 63.00% 0.665 63.42% 0.663 - W2B Bai 63.83% 0.663 63.48% 0.663 63.57% 0.668 63.79% 0.664 - WMB Bai 64.25% 0.661 64.15% 0.679 63.90% 0.677 64.07% 0.679 VIT HBB Bai 64.45% 0.675 64.33% 0.679 63.93% 0.674 64.00% 0.675 VIT WMB Bai 63.42% 0.666 64.87% 0.681 63.62% 0.664 63.59% 0.672 CLB HBB Bai 62.94% 0.657 63.93% 0.678 63.97% 0.678 64.66% 0.687
5 Discussion
5.1 Future work
Firstly, dimensional emotions are also important. In fact, we are labeling MDPE with dimensional emotions (including deception and emotion), which can not only study the impact of dimensional emotional features on lie detection at the individual level, but also explore more emotional clues in the deception process. Secondly, we extracts deep features from some pre-trained models and uses simple models for deception detection. In the future,larger models should be designed to be used for deception detection tasks. Thirdly, this article simply concatenates personality and emotional expression features to assist in deception detection tasks. In the future, more complex feature fusion or model fusion algorithms can be used for deception detection. Finally, in this paper, we only focuses on deception detection tasks, but MDPE includes individual personality, deception, and emotional expression information, which can not only support deception detection tasks, but also tasks such as personality recognition and emotion recognition. Even these individual level information can be used to assist other tasks, as shown in Figure 3. In fact, there have been many studies on the relationship between personality and emotion [61, 62, 63], but the lack of relevant datasets has led to slow research progress in this field, and we believe that MDPE can provide valuable resources for these research directions.
Figure 3: Possible future work on the MDPE
5.2 Limitations
@@
Firstly, although the subjects were required that they must lie about the deception questions, and verified the deceptive questions and content with the Interviewer after the deception experiment, we do not know whether the subjects have actually deceived on the deception questions. Secondly, although our induction videos have been annotated by professionals to demonstrate their reliability and validity, there are still some subject whose feelings are inconsistent with the emotions we expected to induce. This is because different people have different understandings of the video content, triggering different emotions. Thirdly, relying on self-assessment scales for data annotation is a subjective process for subjects, which may lead to bias in subsequent analysis. Different subjects may have significant differences in their perception of emotions. In addition, MDPE only collects native Chinese speakers, there may be cultural differences in deception detection. Finally, gender imbalance among subjects in MPDE is a common issue in human data collection [64, 65].
6 Conclusion
We present MDPE, a dataset for deception detection featuring three categories of data modalities, as well as personality and emotional characteristics. Firstly, diverse modalities carry complementary information that can be jointly exploited. By providing access to multiple synchronized modalities, MDPE enables cross-modal analyses that have the potential to improve the understanding of the relationships between video, audio, and text. Secondly, it can help improve the understanding of deceive behavior, aiming to develop reliable deception detection algorithms and enhance the security issues related to deception in our society. Thirdly, MDPE provides the personality traits and emotional expression characteristics of each subject, which can help analyze the impact of personality and emotional expression on deceive behavior. In addition, MDPE not only supports deception detection models, but also provides conditions for personality recognition and emotion recognition tasks, and can even study the relationship between deception, personality, and emotion, such as using personality features to improve the performance of emotion recognition tasks. Finally, to promote reproducibility, MDPE also provided a set of benchmark experiments. Although the proposed model focuses on deception detection, the use of personality and emotional features also demonstrates the predictive potential of our dataset. They represent a good starting point for future work that researchers and developers can use as a benchmark. By openly sharing MDPE, we hope to ignite new advances in this critical area of affective computing.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See 5.2
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See 5.2
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [Yes]
    2. (b)
      Did you include complete proofs of all theoretical results? [Yes]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See 4.4
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [No]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See 3.2
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See 3.2
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes]

### v2

@@ 4.1. Data Preprocessing
For the visual modality, raw videos are standardized to 30 fps. Faces are cropped and aligned using the DLib Toolkit [1]. Frame-level features are extracted using visual encoders and compressed to video-level via average pooling. For audio, the track is separated using FFmpeg, standardized to 16kHz mono, and acoustic features extracted. For text, transcripts are generated using the Paraformer ASR toolkit [2]. For each sample $x_{i}$, this yields acoustic features $f_{i}^{a}\in\mathbb{R}^{d_{a}}$, textual features $f_{i}^{l}\in\mathbb{R}^{d_{l}}$, and visual features $f_{i}^{v}\in\mathbb{R}^{d_{v}}$, where $\left\{d_{m}\right\}_{m\in\left\{a,l,v\right\}}$ is the feature dimension for each modality.
4.2. Feature Extraction
Feature selection significantly impacts model performance. To guide feature selection, we evaluate distinct feature types under consistent experimental conditions. For visual modality, compared with handcrafted features, deep features extracted from supervised models are useful for facial expression recognition (Li and Deng, 2020). CLIP (Radford et al., 2021) is a multimodal model based on contrastive learning, where training utilizes text and images to construct positive and negative sample pairs. The Vision Transformer (VIT) (Dosovitskiy et al., 2020) is a transformer encoder model, pre-trained in a supervised manner on a large dataset of images. For acoustic modality, the Wav2vec2 (Baevski et al., 2020) masks speech inputs in the latent space and addresses a contrastive task defined on quantized latent representations. It has been widely applied to downstream speech tasks. HUBERT (Hsu et al., 2021) utilizes offline clustering steps to provide aligned target labels for prediction losses. To better distinguish between speakers, a sentence mixture training strategy WavLM (Chen et al., 2022) is proposed, allowing for the unsupervised creation and merging of additional overlapping sentences during the training process. For text modality, we extract the sbert-chinese-general-v2 features, which is based on the bert-base-chinese version of the BERT model. ChatGLM (Du et al., 2021) is an open-source conversational language model, based on the General Language Model (GLM) architecture. It demonstrats exceptional contextual understanding and more efficient inference capabilities. Baichuan (Yang et al., 2023) is an open-source large-scale model with 13 billion parameters. It features a larger size, more extensive training data, and more efficient inference capabilities.
4.3. Model Structure
@@
(1) $h_{i}^{m}=\text{ReLU}\left(f_{i}^{m}W_{m}^{h}+b_{m}^{h}\right),{m\in\left\{a,l,v\right\}}$
(2) $\hat{y}_{i}=\text{softmax}\left(h_{i}^{m}W_{m}^{d}+b_{m}^{d}\right),m\in\left\{a,l,v\right\}$
@@
(3) $h_{i}=\text{Concat}\left(h_{i}^{a},h_{i}^{l},h_{i}^{v}\right)$
(4) $\alpha_{i}=\text{softmax}\left(h_{i}^{T}W_{\alpha}+b_{\alpha}\right)$
@@
Similarly, for personality traits and emotional expression features, feature fusion is achieved through concatenation. Personality features are derived directly from personality scale scores. Emotional expression features, however, are obtained by first training a dedicated emotion recognition model. Subsequently, all emotional expression samples are processed through this model, and the activations from the last fully connected layer are extracted. These activations then undergo average pooling to form the final emotion expression feature vector. It is noted that the model architecture employed for both personality recognition and emotion recognition tasks remains identical.
@@
Table 2. Unimodal results of deception detection. ”P” denotes the addition of personality features and ”E” denotes the addition of emotional features.
Feature Accuracy with P with E with P & E
VIT 60.30% 61.27% 61.43% 61.55%
CLIP-base 58.54% 59.17% 58.32% 59.11%
CLIP-large 57.30% 58.34% 56.97% 57.67%
eGeMAPS 55.86% 57.22% 56.22% 56.89%
HUBERT-base 58.13% 62.38% 59.35% 62.12%
HUBERT-large 60.80% 62.07% 60.34% 61.87%
Wav2vec2-base 58.75% 59.74% 59.99% 59.84%
Wav2vec2-large 60.10% 61.88% 59.32% 62.10%
WavLM-base 61.66% 60.82% 60.16% 60.92%
WavLM-large 57.82% 60.31% 58.02% 60.52%
Sentence-BERT 61.76% 62.34% 63.21% 63.34%
ChatGLM2-6B 60.73% 61.45% 61.45% 61.56%
Baichuan-13B 61.87% 62.90% 63.32% 63.74%
4.4. Implementation Details
The dimension of latent representations is selected from $\left\{64,128,256\right\}$. We employ the Adam optimizer (Kingma and Ba, 2014) with a learning rate chosen from $\left\{10^{-3},10^{-4}\right\}$, a weight decay of $10^{-5}$, and a maximum of 300 training epochs. To mitigate overfitting, dropout (Srivastava et al., 2014) is applied, with rates selected from $\left\{0.2,0.3,0.4,0.5\right\}$. The cross-entropy loss function is used for optimization. For deception detection tasks, each sample comprises 24 answers. We randomly select 5 answers per sample (3 truthful, 2 deceptive) for validation, reserving the remaining 19 for training. All experiments are repeated five times with randomized initializations, and results report average performance to ensure statistical reliability. For personality and emotion recognition tasks, the dataset is split into 133 training samples, 40 validation samples, and 40 test samples; these tasks utilize root mean square error (RMSE) as the loss function.
For deception detection, accuracy was selected as the evaluation metric. For personality recognition, we employed the mean accuracy (A), defined as follows:
(5) $A=1-\frac{1}{N^{t}}\sum_{i}^{N^{t}}\left|\mathbf{Y}_{i}^{P}-\mathbf{P}_{i}\right|$
This metric is widely adopted in personality recognition tasks (Ponce-López et al., 2016; Zhang et al., 2019). For emotion recognition, the root mean square error (RMSE) was used.
Table 3. Multimodal results of deception detection.
V A T Accuracy with P with E with P & E
VIT HBB - 61.76% 61.53% 61.14% 62.68%
VIT WMB - 60.88% 60.22% 60.02% 60.72%
CLB HBB - 61.02% 60.14% 60.53% 61.45%
VIT - Bai 63.31% 63.47% 63.31% 63.48%
CLB - Bai 63.31% 63.52% 63.38% 63.10%
CLL - Bai 64.15% 63.94% 63.98% 64.04%
- HBL Bai 62.69% 63.42% 63.00% 63.42%
- W2B Bai 63.83% 63.48% 63.57% 63.79%
- WMB Bai 64.25% 64.15% 63.90% 64.07%
VIT HBB Bai 64.45% 64.33% 63.93% 64.00%
VIT WMB Bai 63.42% 64.87% 63.62% 63.59%
CLB HBB Bai 62.94% 63.93% 63.97% 64.66%
4.5. Experiment Results
This section establishes the benchmark for MDPE, designed to guide feature selection and inform the development of robust feature extractors.
Unimodal deception detection results (Table 2) indicate that the textual modality significantly outperforms visual and acoustic modalities. This suggests that textual cues in our dataset provide more salient deceptive indicators. Subsequent multimodal fusion experiments (Table 3) reveal that integrating complementary modalities consistently enhances performance. This improvement aligns with the premise that deception manifests across multiple channels, enabling models to better comprehend video content and detect deception more accurately. Notably, while most unimodal features benefit from fusion, combining visual and acoustic features yields negligible gains or performance degradation. This implies that textual features stabilize model performance, a finding consistent with human deception judgment, where content-based (textual) cues typically dominate over visual or acoustic signals. Further exploration of deceptive indicators within visual and acoustic modalities is warranted.
Incorporating personality features consistently improves deception detection performance, underscoring their importance for the task. While emotion features also enhance performance, their contribution is less pronounced than personality features and occasionally detrimental. This discrepancy may arise because personality traits serve as direct indicators, whereas emotion features depend on the quality of upstream emotion recognition models. Future work should explore more effective methods for leveraging emotional expression features. Combining both personality and emotion features achieves the highest unimodal deception detection performance, confirming their relevance and demonstrating the viability of individual-difference-based modeling.
For personality recognition (Table 3), acoustic features surpass visual/textual features in isolation, highlighting the predictive value of vocal cues. Full multimodal fusion achieves optimal performance, indicating cross-modal complementarity. In emotion recognition, textual features provide the strongest cues, and multimodal integration significantly outperforms unimodal approaches, emphasizing the necessity of cross-modal fusion for robust emotion understanding.
Table 4. Experiment results of personality recognition and emotion recognition.
Feature Personality Emotion
Val Test Val Test
VIT 90.70% 92.2% 1.288 1.351
ClipVIT-B16 92.06% 91.75% 1.272 1.333
ClipVIT-L14 91.81% 91.47% 1.277 1.339
HUBERT-base 92.01% 92.65% 1.286 1.330
HUBERT-large 91.62% 92.66% 1.290 1.336
Wav2vec2-base 92.27% 93.16% 1.281 1.337
Wav2vec2-large 92.32% 92.96% 1.287 1.342
WavLM-base 91.64% 92.53% 1.280 1.336
WavLM-large 91.82% 92.55% 1.287 1.344
Sentence-BERT 92.23% 93.04% 1.274 1.327
ChatGLM2-6B 91.75% 92.63% 1.276 1.334
Baichuan-13B 92.30% 93.11% 1.274 1.332
VIT W2V - 90.71% 92.36% 1.272 1.334
- W2V BAI 92.34% 93.35% 1.273 1.332
VIT - BAI 90.74% 92.31% 1.275 1.334
VIT W2V BAI 92.16% 93.43% 1.270 1.229
5. Limitations
@@
Firstly, although the subjects were required that they must lie about the deception questions, and verified the deceptive questions and content with the Interviewer after the deception experiment, we do not know whether the subjects have actually deceived on the deception questions. Secondly, relying on self-assessment scales for data annotation is a subjective process for subjects, which may lead to bias in subsequent analysis. Different subjects may have significant differences in their perception of emotions. In addition, MDPE only collects native Chinese speakers, there may be cultural differences in deception detection. Finally, gender imbalance among subjects in MPDE is a common issue in human data collection (D’Mello et al., 2022; Pinho-Gomes et al., 2022).
6. Conclusion
We introduce the Multimodal Deception Detection Dataset (MDPE), comprising video, audio, and textual modalities, supplemented with personality and emotion annotations. This dataset enables cross-modal analysis to investigate complementary relationships between modalities, thereby advancing robust deception detection methodologies relevant to societal security. Furthermore, MDPE facilitates research into the influence of personality traits and emotional states on deceptive behavior. Beyond its primary application, the dataset supports auxiliary tasks such as personality and emotion recognition, as well as joint analyses of deception-personality-emotion interactions. Benchmark experiments are provided to ensure reproducibility and establish a foundation for future work. By publicly releasing MDPE, we aim to stimulate progress in this critical area of affective computing.